OpenStudy (anonymous):

Write the function whose graph is the graph of y=(x+1)^2 but is reflected about the x-axis.

OpenStudy (anonymous):

stick a minus sign in front of it

OpenStudy (anonymous):

y=-(x+1)^2 ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@laurenn_marie7

OpenStudy (anonymous):

That's right laurenn_marie7. Whenever you have to reflect a function in the x-axis, you stick a minus sign in front of it. So y = (x + 1)^2 becomes y = -(x + 1)^2

OpenStudy (anonymous):

Thank you @genius12 !! :) would y=x^2 be y=(1/7x)^2 when it is horizontally stretched by 7?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you!!! @genius12 last one sorry, this one asks what y= sq root of x would be after being shifted down 5 units, reflected about the x-axis and reflected about the y-axis? I don't think mine is right... I got -(sq.root of -x)+5 ?

OpenStudy (anonymous):

\[\sqrt{x} ... I got -\sqrt{-x}+5\]